OpenStudy (anonymous):

xy (x+y)=3 x^3+y^3=8 ---> (x+y)^3 =?

OpenStudy (asnaseer):

try expanding \((x+y)^3\) - you should get a clue from that

OpenStudy (anonymous):

i try but i dont understand

OpenStudy (asnaseer):

what do you get when you expand that?

OpenStudy (anonymous):

= (x+y) ( x^2 - xy + y^2)

OpenStudy (asnaseer):

that is not correct - there should not be a mius sign there

OpenStudy (asnaseer):

\[(x+y)^3=(x+y)(x+y)(x+y)=(x+y)(...)\]

OpenStudy (asnaseer):

you need to replace the ... with the expansion of this: \((x+y)(x+y)=?\)

OpenStudy (asnaseer):

you can break this up as follows:\[(x+y)(x+y)=x(x+y)+y(x+y)=?\]

OpenStudy (anonymous):

I am not understanding

OpenStudy (asnaseer):

expand the last expression I typed above, i.e.:\[(x+y)(x+y)=x(x+y)+y(x+y)=?\]

OpenStudy (asnaseer):

this comes from: